OpenStudy (anonymous):

[2.05] In a certain county, the number of Charter Schools is eleven less than twice the number of Alternative Schools. There are 35 Charter Schools in the county. Which of the following equations would you use to find the number of Alternative Schools, a, in the county? (5 points) 2a − 11 = 35 2a = 35 − 11 2(a − 11) = 35

OpenStudy (compassionate):

c If the original equation says: 2a-11, then you will want to subtract the charter schools from the alternative times two.

OpenStudy (aripotta):

i don't think it's c

OpenStudy (aripotta):

you even said yourself, it's 2a - 11 = 35

OpenStudy (compassionate):

C Schools 11 less than 2 * A schools, so 2a - 11 = c. But c = 35, so 2a - 11 = 35

OpenStudy (anonymous):

a-11 ---- 35? 2

OpenStudy (anonymous):

=35*

OpenStudy (aripotta):

oh ok. i see what you mean

OpenStudy (aripotta):

it's the first option

OpenStudy (anonymous):

Thanks guys.